NO_{3} STATES ACCESSED BY PHOTODETACHMENT AND DARK-STATE SPECTROSCOPY: WHAT DO WE KNOW?

\underline{JOHN F. STANTON}, MITCHIO OKUMURA, Institute for Theoretical Chemistry, Department of Chemistry, University of Texas at Austin, Austin, TX 78712; Arthur Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena CA 91125.

A good argument can be made that the atmospherically relevant NO_{3} radical is the most complicated tetraatomic molecule. Five electronic states (the ground X̃^{2}A′_{2} state, and the two degenerate Ã^{2}E′′ and B̃^{2}E′ states) lie within 2 eV, and are coupled by several strong interactions. Amongst these is a strong linear and quadratic Jahn-Teller effect in the Ã^{2}E′′ state, and a profound pseudo-Jahn-Teller interaction which couples the X̃^{2}A′_{2} ground state with the second excited state. The latter is almost strong enough to break the D_{3h} equilibrium symmetry of the ground state, and severely perturbs vibrational levels associated with the two in-plane modes of e′ symmetry. This talk will report ab initio calculations based on the equation-of-motion coupled cluster model, its application to parametrize various vibronic model Hamiltonians, and simulations of the photodetachment spectrum of NO_{3}^{−} as well as the recent experiments of Okumura and coworkers in which the (dipole forbidden) Ã^{2}E′′ state is accessed by direct absorption from the ground state.